OpenStudy (anonymous):

Find the volume of a solid by rotating the region bounded by the given curves about the specified line: xy=1, y=0, x=1, x=2; about x=-1

OpenStudy (anonymous):

I think method of shells would work here, integrating dx as thickness of shell.

OpenStudy (anonymous):

I got that far, but I don't know what to set as the radii

OpenStudy (anonymous):

Since the axis of rotation is at x=-1, the radius will be 1+x. Drawing a quick sketch of the curve and the boundaries makes it easier to see.

OpenStudy (anonymous):

The generic differential volume would be dV=2π*r*y*dx r=(1+x), y = 1/x. Integrate between 1 and 2 and that should do it. I'm a little rusty at these, so I might be off somewhere.